# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

### v1

@@ Abstract
Large vision-language models (LVLMs) have witnessed significant progress on visual understanding tasks. However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation. To tackle this issue, we first identify the drawbacks of existing solutions (i.e., insufficient and irrelevant visual descriptions, and limited multi-modal capacities). We then decompose visual reasoning process into two stages: visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason. This framework features multi-run proactive perception and decoupled vision-reasoning capabilities. Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions. Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs. Our extensive experiments demonstrate that ProReason outperforms both existing multi-step reasoning frameworks and passive peer methods on a wide range of benchmarks for both open-source and closed-source models. In addition, with the assistance of LLMs, ProReason achieves a performance improvement of up to 15% on MMMU benchmark. Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones.
@@ 1 Introduction
In recent years, large language models (LLMs) (Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024). This progress has also sparked interest in developing large vision-language models (LVLMs) (Chen et al., 2024a; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks. However, state-of-the-art (SOTA) LVLMs still struggle to integrate visual understanding with textual reasoning simultaneously due to inherent modality differences and deficient training data. For example, Ghosh et al. (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks such as visual chart understanding and math reasoning, resulting in performance degradation. Figure 2 illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
@@
To address both shortcomings, we propose a novel multi-step multi-modal reasoning framework named ProReason, which features proactive (i.e., question-oriented and reasoning-involved) visual extraction and disentangled vision-reasoning capabilities. Specifically, we first decouple multi-modal reasoning capacity into two sub-tasks: visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom). The former focuses on understanding images and converting visual details into textual form, while the latter integrates the gathered information to draw final conclusions. As illustrated in Figure 1, our framework can be divided into three stages: Action, Judgment, and Summary, with each involving different agents. In the Action stage, a Dispatcher selectively engages a Vision Expert to capture additional visual information, or a Reasoning Expert to derive more insights. Unlike passive methods, all three agents operate based on the given question and current descriptions, effectively avoiding information redundancy or insufficiency. Subsequently, a Referee in the Judgment stage evaluates the sufficiency of known information, deciding whether to return to the Action stage or proceed to the Summary stage, where all information is consolidated for a Summarizer to generate the final answer. Notably, the disentangled vision-reasoning does not require vision-irrelevant roles (i.e., Reasoning Expert, Referee, and Summarizer) to be performed by LVLMs, enabling the seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby addressing the limitations of LVLMs.
Empirically, we evaluate ProReason across four challenging visual reasoning benchmarks with both open-source and closed-source models. Extensive experiments demonstrate that ProReason achieves significant and consistent performance improvement over both existing reasoning frameworks and state-of-the-art passive methods across all the benchmarks, with a peak enhancement as 13.2%. Besides, our comparative analysis highlights the necessity and relative importance of each sub-agent, as well as the advantages of proactive information acquisition. With the verified superiority over the simultaneous inherent usage of both capabilities, the decoupled perspective not only demonstrates that textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, but also exemplifies the feasibility of LLM-assisted LVLM reasoning, exhibiting performance improvements of up to 15% on the MMMU benchmarks (Yue et al., 2023).
@@
  1. We propose a novel multi-modal reasoning framework named ProReason, featuring iterative proactive perception and decoupled vision-reasoning capabilities.
  2. Empirically, we identify the drawbacks of existing solutions (i.e., insufficient and irrelevant visual descriptions, and limited multi-modal capacities), and validate the successful mitigation of these issues with ProReason by the superior performance over peer methods.
@@ 2 Preliminary Observations
We first analyze the behaviors of several existing methods on the MMMU (Yue et al., 2023) dataset, a challenging multi-modal benchmark requiring comprehensive college-level knowledge and fine-grained reasoning abilities. All experiments are conducted with three recent LVLMs for robustness: Llama3-LLaVA-NeXT-8B (Li et al., 2024a), LLaVA-OneVision-Qwen2-7B-OV (Li et al., 2024b) and Qwen-VL-Chat (Bai et al., 2023). Further details are provided in Sec. 4.1.

### v2

@@ Abstract
Large vision-language models (LVLMs) have witnessed significant progress on visual understanding tasks. However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation. To tackle this issue, we first identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions). We then decompose visual reasoning process into two stages: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason. This framework features decoupled vision-reasoning capabilities and multi-run proactive perception. Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions. Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs. Our extensive experiments demonstrate that ProReason outperforms both existing multi-step reasoning frameworks and passive peer methods on various benchmarks for both open-source and closed-source models. More impressively, with the assistance of LLMs, ProReason achieves a performance improvement of up to 15% on MMMU benchmark. Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones [^1] Code is available at https://github.com/lian-tian-mo-zun/Pro_Reason.
ProReason: Multi-Modal Proactive Reasoning
with Decoupled Eyesight and Wisdom
Jingqi Zhou, Sheng Wang, Jingwei Dong, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
@@ 1 Introduction
In recent years, large language models (LLMs) (Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024). This progress has also sparked interest in developing large vision-language models (LVLMs) (Chen et al., 2024a; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks. However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences. For example, Ghosh et al. (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks such as visual chart understanding and math reasoning, resulting in performance degradation. Figure 2.b illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
@@
To address both shortcomings, we propose a novel multi-step multi-modal reasoning framework named ProReason, which features disentangled vision-reasoning capabilities and proactive (i.e., question-oriented and reasoning-involved) visual extraction. As illustrated in Figure 1, we first decouple multi-modal reasoning capacity into two sub-tasks: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), based on the capacity of LVLMs. The former focuses on understanding images and converting visual details into textual form, while the latter integrates the gathered information to draw final conclusions, with each involving different agents. In the proactive visual perception stage, a Dispatcher decomposes the original question into visual or reasoning-related sub-questions and selectively engages a Vision Expert to capture additional visual information, or a Reasoning Expert to derive extra insights. Subsequently, a Referee evaluates the sufficiency of known information, deciding whether to continue the proactive visual perception stage or proceed to the textual reasoning stage, where all information is consolidated for a Summarizer to generate the final answer. Unlike passive methods, all sub-agents operate based on the given question and current descriptions, effectively avoiding information redundancy or insufficiency. Notably, the disentangled vision-reasoning does not require vision-irrelevant roles (i.e., Reasoning Expert, Referee, and Summarizer) to be performed by LVLMs, enabling the seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby alleviate the limitations of LVLMs.
Empirically, we evaluate ProReason across four challenging visual reasoning benchmarks with both open-source and closed-source models. Extensive experiments demonstrate that ProReason achieves significant and consistent performance improvement over both existing reasoning frameworks and state-of-the-art passive methods across all the benchmarks, with a peak enhancement as 13.2% (Sec. 4.2). Besides, our comparative analysis highlights the necessity and relative importance of each sub-agent (Sec. 4.3), as well as the crucial role of decoupling (Sec. 4.4). With the verified superiority over the simultaneously inherent usage of both capabilities, the decoupled perspective not only demonstrates that textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks (Sec. A.3.1), but also exemplifies the feasibility of LLM-assisted LVLM reasoning, exhibiting performance improvements of up to 15% on the MMMU benchmarks (Yue et al., 2023).
@@
  1. We propose a novel multi-modal reasoning framework named ProReason, featuring decoupled vision-reasoning and iterative proactive perception capabilities.
  2. Empirically, we identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions), and validate the successful mitigation of these issues with ProReason by the superior performance over peer methods.
@@ 2 Preliminary Observations
We first analyze the behaviors of several existing methods on the MMMU (Yue et al., 2023) dataset, a challenging multi-modal benchmark requiring comprehensive college-level knowledge and fine-grained reasoning abilities. All experiments are conducted with three recent LVLMs for the robustness of the results: Llama3-LLaVA-NeXT-8B (Li et al., 2024a), LLaVA-OneVision-Qwen2-7B-OV (Li et al., 2024b) and Qwen-VL-Chat (Bai et al., 2023). Further details are provided in Sec. 4.1.